# Scalable Hypergraph Structure Learning with Diverse Smoothness Priors

### v1

@@ V-A Co-authorship Networks
We experiment on two real world co-authorship networks: DBLP and Cora [29]. For each dataset, nodes are considered authors, and a hyperedge groups nodes if the nodes’ corresponding authors have written a paper together. We use a subset of the overall datasets to experiment on, and these subsets are the same that Pena-Pena et al. [23] used. Therefore, we can compare our learning method and smoothness terms to their tensor based method PDL-HGSP. We also compare results with the HGSI method from Tang et al. [22] and the K-NN approach in the node feature space from Gao et al. [17]. The experiment is performed on both a uniform and non-uniform cardinality hypergraph from both DBLP and Cora. In the uniform scenarios, the set of cardinalities $\mathcal{K}=\{3\}$. In the non-uniform scenarios, $\mathcal{K}=\{2,3\}$. Table II details the structural parameters of the ground truth hypergraphs used.
TABLE II: Number of nodes and hyperedges for experimental co-authorship networks.
@@
Figure 4: Bar graphs of overlap percentage between $E_{truth}(\mathcal{H})$ and $E_{KNN}(\mathcal{H})$. Red bars indicate the percentage is less than $100\%$. Green bars indicate the overlap percentage has reached $100\%$. The x-axis is labeled in terms of the number of nearest neighbors used dependent on $k_{l}$, the cardinalities from set $\mathcal{K}$. Each bar plot corresponds to one of the co-authorship networkst: (a) Uniform DBLP, (b) Non-Uniform DBLP, (c) Uniform Cora, and (d) Non-Uniform Cora.
Now, with regard to signals on the DBLP and Cora datasets, there is no given set of time-series observations. Since our method is based upon a smoothness prior, we first need to generate smooth signals across the network structures. We adopt the same process used by Tang et al. who uses a multivariate Gaussian distribution for smooth hypergraph signals, similar to processes used in graph signal processing [8]. Using the Laplacian matrix of the hypergraph star expansion $\mathbf{L}_{s}$, the signal distribution can be modeled as
@@
where $\mathbf{L}_{s}^{\dagger}$ is the pseudo-inverse of $\mathbf{L}_{s}$, $\mathbf{0}$ is a vector of all zeros, $\mathbf{x}_{\mathbf{V}}\in\mathbb{R}^{N}$ is the signal across hypergraph nodes, and $\mathbf{x}_{\mathcal{E}}\in\mathbb{R}^{Y}$ is the signal across the hyperedges. After sampling from the multivariate Gaussian distribution, we only kept $\mathbf{x}_{\mathbf{V}}$ as one hypergraph signal observation, and we stored all hypergraph signal observations in $\mathbf{X}$. We also add some noise along the diagonal of $\mathbf{L}_{s}$ on the order of $1\times 10^{-3}$. We note that, in this context, $\mathbf{L}_{s}$ is the Laplacian of the star expansion of the ground truth hypergraphs from both the DBLP and Cora datasets and is a known quantity. We used $P=250$ hypergraph signal observations to produce $\mathbf{X}$ which was kept constant across all learning methods per scenario.
@@
We started with an analysis of the overlap between the K-NN selected hyperedges and the ground truth for all co-authorship networks. Using $\mathcal{K}$, we selected the $k_{l}-1$ nearest neighbors of each node from $\mathbf{X}$ to form hyperedges. We then computed the overlap percentage between the set of ground truth hyperedges and the K-NN hyperedges. By this we mean that, for each hyperedge in the ground truth set $E_{truth}(\mathcal{H})$, we determined if it was present in the K-NN hyperedge set $E_{KNN}(\mathcal{H})$. If $E_{truth}(\mathcal{H})\subseteq E_{KNN}(\mathcal{H})$ then this is a $100\%$ overlap rate. We then increased the number of nearest neighbors by one and recomputed the overlap percentage for the new, larger K-NN set. This was repeated for a steadily increasing count of nearest neighbors. The computation of $E_{KNN}(\mathcal{H})$ was performed using Algorithm 1. The results are shown in the bar graphs of Fig. 4.
@@
From the results of Fig. 4, we note that in the uniform hypergraph scenarios, the overlap percentage starts at a higher threshold than that of the non-uniform cases. This can be attributed to two main factors. The first is that the uniform scenarios have less $|E(\mathcal{H})|$ than the non-uniform scenarios. The second is that a mixed cardinality hypergraph has a much larger set of possible hyperedges making it less straightforward to select valid hyperedges. In the uniform cases, it takes significantly less neighbors in the feature space to achieve $E_{truth}(\mathcal{H})\subseteq E_{KNN}(\mathcal{H})$ than that of the non-uniform cases. In fact, the minimum number of neighbors for this in the uniform and non-uniform scenarios is $k_{l}$ and $k_{l}+3$, respectively. The use of $k_{l}-1$ nearest neighbors always falls below 100% overlap, demonstrating the limitation of the method of HGSI since, after the reduction of hyperedge possibilities, it is impossible for it to reproduce a hypergraph that completely corresponds to the ground truth. We also find that there is indeed a reasonable point where 100% overlap can be achieved, therefore validating our choice to incorporate a K-NN hyperedge reduction method with number of neighbors greater than $k_{l}-1$. In the rest of the co-authorship network experiments, we used the corresponding number of neighbors that produced 100% overlap in Fig. 4 to reduce the size of $\mathbf{w}$ and, consequently, $\mathbf{z}$.
@@
We used $\mathbf{z}$, calculated from $\mathbf{X}$ and the reduced hyperedge set, along with Algorithm 2 to get the resultant $\mathbf{w}$. To refine our choice of $\alpha$ and $\beta$, we used a hyperparameter grid search on a log scale. The statistical results of recovering hypergraph structures for both datasets in the uniform and non-uniform scenarios can be found in Table III. A visual result of the DBLP uniform testing using HSLS can be found in Fig. 5.
TABLE III: Co-authorship network experimental metric results.
@@
DBLP PDL-HGSP 0.9130 0.8750 0.9545 0.7143 0.5932 0.8974
Gao et al. 0.8000 0.7143 0.9091 0.5625 0.4045 0.9231
HGSI 0.9091 0.9091 0.9091 0.4103 0.4103 0.4103
@@
Cora PDL-HGSP 1.0000 1.0000 1.0000 0.8000 0.7755 0.8261
Gao et al. 0.7200 0.6429 0.8182 0.5755 0.4396 0.8333
HGSI 0.8182 0.8182 0.8182 0.4583 0.4583 0.4583
@@
HSLS with the other selected total variation terms performed decent in the uniform cases, often producing better results than the compared methods. However, only the Max-Absolute total variation maintained some of this statistical response in the non-uniform case, which was a foreseen outcome discussed in Sec. II. Comparatively, HSLS with Max-Square still outperformed the other selected total variations. Since F1-score is the combination of precision and recall, it is the most apt description of overall performance. Therefore, our method was the strongest performer across the board when combined with Max-Square. We also note that all tested methods demonstrate a steep decrease in statistical output when used in a non-uniform structure application, thus demonstrating the more challenging nature of increasingly complex networks.
@@ V-B Lung Cancer Mortality Rates
The goal is to learn a uniform hypergraph that captures a smooth structure across the nodes to evaluate the emphasis our method puts on hypergraph smoothness. To do this, we first create a baseline hypergraph using K-NN on the mortality data to create a uniform cardinality hypergraph where $\mathcal{K}=\{3\}$. K-NN is biased towards creating hyperedges that are locally smooth, but not necessarily globally across the entire hypergraph. However, as we previously investigated with the co-authorship networks, K-NN selected hyperedges can produce reasonable approximations of globally smooth hypergraphs. To that end, we used each of the four proposed total variations on the baseline K-NN hypergraph structure. Then, we used HSLS to recover hypergraphs with the mortality data. We then calculated the corresponding total variation of our learned hypergraphs for comparison with the baseline. The idea here is that the baseline hypergraph should be relatively smooth, due to K-NN local properties, but our method should encourage better results that are globally smooth. Since the magnitude of the total variation calculations are dependent on the number of hyperedges, we constrain our optimization to have at least as many hyperedges as the baseline.
@@
To further demonstrate the experimental smoothness properties, we introduced two other datasets. The first contains smoking rates [31] and the second contains temperature readings [32]. Both datasets are reported on a year by year basis from 2011-2019 and are therefore of a similar format to the mortality data. We then used the hypergraph structure learned on the mortality data to calculate the corresponding total variation with the smoking and temperature readings. It is well known that the onset of lung cancer and smoking are certainly correlated factors [33]. Furthermore, the correlation between lung cancer and temperature is at best ambiguous, but certainly not as strong as smoking. To demonstrate this idea, Fig. 6 contains color maps of the U.S. based on average lung cancer mortality (Fig. 6 left), average smoking rates (Fig. 6 middle), and average temperature (Fig. 6 right) across the year. The data has been normalized so the figures are on comparable scales. Visually, it is clear that the coloring of the mortality and smoking maps share similar trends, especially in the southeast and northeast regions. These trends become less similar when comparing these two maps to the temperature map which gradients across the U.S. from north to south. So, we expect that if our method can successfully learn a smooth hypergraph on the mortality data, then correlated factors, such as smoking rates, would also be approximately smooth on the same structure. We also expect that factors uncorrelated to the mortality data, such as temperature, should be less smooth across the structure than that of the correlated factors. The results of using K-NN, the learned hypergraphs, and correlated factors are all reported in Table IV.
@@
TABLE IV: Results of uniform hypergraph learning using the structure recovered from mortality data.
@@
K-NN Total Variation
59.3 148.5 63.0 28.5
K-NN $|\mathcal{E}|$ 32 32 32 32
HSLS Total Variation
41.9 142.7 60.4 11.5
HSLS $|\mathcal{E}|$
38 32 32 34
Smoking Total Variation
305.5 431.1 206.1 162.0
Temperature Total Variation
2006.7 922.2 500.3 1184.6

### v2

@@ V-A Co-authorship Networks
We experiment on two real world co-authorship networks: DBLP and Cora [48]. For each dataset, nodes are considered authors, and a hyperedge groups nodes if the nodes’ corresponding authors have written a paper together. We use a subset of the overall datasets to experiment on, and these subsets are the same that Pena-Pena et al. [26] used. Therefore, we can compare our learning method and smoothness terms to their tensor based method PDL-HGSP. We also compare results with the HGSI method from Tang et al. [38] and the K-NN approach in the node feature space from Gao et al. [28]. The experiment is performed on both a uniform and non-uniform cardinality hypergraph from both DBLP and Cora. In the uniform scenarios, the set of cardinalities $\mathcal{K}=\{3\}$. In the non-uniform scenarios, $\mathcal{K}=\{2,3\}$. Table I details the structural parameters of the ground truth hypergraphs used.
TABLE I: Number of nodes and hyperedges for experimental co-authorship networks.
@@
Figure 4: Bar graphs of overlap percentage between $E_{truth}(\mathcal{H})$ and $E_{KNN}(\mathcal{H})$. Red bars indicate the percentage is less than $100\%$. Green bars indicate the overlap percentage has reached $100\%$. The x-axis is labeled in terms of the number of nearest neighbors used dependent on $k_{l}$, the cardinalities from set $\mathcal{K}$. Each bar plot corresponds to one of the co-authorship networks: (a) Uniform DBLP, (b) Non-Uniform DBLP, (c) Uniform Cora, and (d) Non-Uniform Cora.
Now, with regard to signals on the DBLP and Cora datasets, there is no given set of time-series observations. Since our method is based upon a smoothness prior, we first need to generate smooth signals across the network structures. We adopt the same process introduced by Tang et al. [38] who use a multivariate Gaussian distribution for smooth hypergraph signals, similar to processes used in graph signal processing [15]. Using the Laplacian matrix of the hypergraph star expansion $\mathbf{L}_{s}$, the signal distribution can be modeled as
@@
where $\mathbf{L}_{s}^{\dagger}$ is the pseudo-inverse of $\mathbf{L}_{s}$, $\mathbf{0}$ is a vector of all zeros, $\mathbf{x}_{\mathbf{V}}\in\mathbb{R}^{N}$ is the signal across hypergraph nodes, and $\mathbf{x}_{\mathcal{E}}\in\mathbb{R}^{Y}$ is the signal across the hyperedges. After sampling from the multivariate Gaussian distribution, we only kept $\mathbf{x}_{\mathbf{V}}$ as one hypergraph signal observation, making this generative process resemble marginalization over a Gaussian Markov Random Field induced through sampling. We stored all hypergraph signal observations in $\mathbf{X}$. We also add some noise along the diagonal of $\mathbf{L}_{s}$ on the order of $1\times 10^{-3}$. We note that, in this context, $\mathbf{L}_{s}$ is the Laplacian of the star expansion of the ground truth hypergraphs from both the DBLP and Cora datasets and is a known quantity. We used $P=250$ hypergraph signal observations to produce $\mathbf{X}$ which was kept constant across all learning methods per scenario.
@@
We started with an analysis of the overlap between the K-NN selected hyperedges and the ground truth for all co-authorship networks. Using $\mathcal{K}$, we selected the $k_{l}-1$ nearest neighbors of each node from $\mathbf{X}$ to form hyperedges. We then computed the overlap percentage between the set of ground truth hyperedges and the K-NN hyperedges. By this we mean that, for each hyperedge in the ground truth set $E_{truth}(\mathcal{H})$, we determined if it was present in the K-NN hyperedge set $E_{KNN}(\mathcal{H})$. If $E_{truth}(\mathcal{H})\subseteq E_{KNN}(\mathcal{H})$ then this is a $100\%$ overlap rate. We then increased the number of nearest neighbors by one and recomputed the overlap percentage for the new, larger K-NN set. This was repeated for a steadily increasing count of nearest neighbors. The computation of $E_{KNN}(\mathcal{H})$ was performed using Algorithm 1, where $\mathcal{K}=\{3\}$ and $\mathcal{K}=\{3,4\}$ in the uniform and non-uniform scenarios, respectively, and $\mathcal{R}$ was continuously updated with the increasing nearest neighbor count starting at $k_{l}-1$. The results are shown in the bar graphs of Fig. 4.
@@
From the results of Fig. 4, we note that in the uniform hypergraph scenarios, the overlap percentage starts at a higher threshold than that of the non-uniform cases. This can be attributed to two main factors. The first is that the uniform scenarios have less $|E(\mathcal{H})|$ than the non-uniform scenarios. The second is that a mixed cardinality hypergraph has a much larger set of possible hyperedges making it less straightforward to select valid hyperedges. In the uniform cases, it takes significantly less neighbors in the feature space to achieve $E_{truth}(\mathcal{H})\subseteq E_{KNN}(\mathcal{H})$ than that of the non-uniform cases. In fact, the minimum number of neighbors for this in the uniform and non-uniform scenarios is $k_{l}$ and $k_{l}+3$, respectively. The use of $k_{l}-1$ nearest neighbors always falls below 100% overlap, demonstrating the limitation of the method of HGSI since, after the reduction of hyperedge possibilities, it is impossible for it to reproduce a hypergraph that completely corresponds to the ground truth. We also find that there is indeed a reasonable point where 100% overlap can be achieved, therefore validating our choice to incorporate a K-NN hyperedge reduction method with number of neighbors greater than $k_{l}-1$. In the rest of the co-authorship network experiments, we used =$\mathcal{R}=\{3\}$ in the uniform and =$\mathcal{R}=\{5,6\}$ in the non-uniform pre-processing with Algorithm 1. This corresponds to the number of neighbors that produced 100% overlap in Fig. 4.
@@
We used $\mathbf{z}$, calculated from $\mathbf{X}$ and the reduced hyperedge set $\hat{E}(\mathcal{H})$, along with Algorithm 2 to get the resultant $\mathbf{w}$. To refine our choice of $\alpha$ and $\beta$, we used a hyperparameter grid search on a log scale. The statistical results of recovering hypergraph structures for both datasets in the uniform and non-uniform scenarios can be found in Table II. A visual result of the DBLP uniform testing using HSLS can be found in Fig. 5.
TABLE II: Co-authorship network experimental metric results.
@@
DBLP PDL-HGSP [26] 0.9130 0.8750 0.9545 0.7143 0.5932 0.8974
Gao et al. [28] 0.8000 0.7143 0.9091 0.5625 0.4045 0.9231
HGSI [38] 0.9091 0.9091 0.9091 0.4103 0.4103 0.4103
@@
Cora PDL-HGSP [26] 1.0000 1.0000 1.0000 0.8000 0.7755 0.8261
Gao et al. [28] 0.7200 0.6429 0.8182 0.5755 0.4396 0.8333
HGSI [38] 0.8182 0.8182 0.8182 0.4583 0.4583 0.4583
@@
HSLS with the other selected total variation terms besides Max-Square performed decent in the uniform cases, often producing better results than the compared methods. However, only the Max-Absolute total variation maintained some of this statistical response in the non-uniform case. Comparatively, HSLS with Max-Square still outperformed the other selected total variations. Since F1-score is the combination of precision and recall, it is the most apt description of overall performance. Therefore, our method was the strongest performer across the board when combined with Max-Square. We also note that all tested methods demonstrate a steep decrease in statistical output in a non-uniform structure application, thus demonstrating the more challenging nature of increasingly complex networks.
@@ V-B Lung Cancer Mortality Rates
The goal was to learn a uniform hypergraph that captures a smooth structure across the nodes to evaluate the emphasis our method puts on hypergraph smoothness. To do this, we first created a baseline hypergraph using K-NN on the mortality data to generate a uniform cardinality hypergraph where $\mathcal{K}=\{3\}$. K-NN is biased towards creating hyperedges that are locally smooth, but not necessarily globally across the entire hypergraph. However, as we previously investigated with the co-authorship networks, K-NN selected hyperedges can produce reasonable approximations of globally smooth hypergraphs. To that end, we applied each of the four proposed total variations to the baseline K-NN hypergraph structure. Then, we used HSLS to recover hypergraphs with the mortality data. We then calculated the corresponding total variation of our learned hypergraphs for comparison with the baseline. The idea here is that the baseline hypergraph should be relatively smooth, due to K-NN local properties, but our method should encourage better results that are globally smooth. Since the magnitude of the total variation calculations are dependent on the number of hyperedges, we constrain our optimization to have at least as many hyperedges as the baseline.
@@
To further demonstrate the experimental smoothness properties, we introduced two other datasets. The first contains smoking rates [50] and the second contains temperature readings [51]. Both datasets are reported on a year by year basis from 2011-2019 and are therefore of a similar format to the mortality data. We then used the hypergraph structure learned on the mortality data to calculate the corresponding total variation with the smoking and temperature readings. It is well known that the onset of lung cancer and smoking are certainly correlated factors [52]. Furthermore, the correlation between lung cancer and temperature is at best ambiguous, but certainly not as strong as smoking. To demonstrate this idea, Fig. 6 contains color maps of the U.S. based on average lung cancer mortality (Fig. 6 left), average smoking rates (Fig. 6 middle), and average temperature (Fig. 6 right) across the year. The data has been normalized so the figures are on comparable scales. Visually, it is clear that the coloring of the mortality and smoking maps share similar trends, especially in the southeast and northeast regions. These trends become less similar when comparing these two maps to the temperature map which gradients across the U.S. from north to south. So, we expect that if our method can successfully learn a smooth hypergraph on the mortality data, then correlated factors, such as smoking rates, would also be approximately smooth on the same structure. We also expect that factors uncorrelated to the mortality data, such as temperature, should be less smooth across the structure than that of the correlated factors. The results of using K-NN, the learned hypergraphs, and correlated factors are all reported in Table III.
@@
TABLE III: Results of uniform hypergraph learning using the structure recovered from mortality data.
@@
K-NN TV 59.3 148.5 63.0 28.5
K-NN $|E(\mathcal{H})|$ 32 32 32 32
HSLS TV 41.9 142.7 60.4 11.5
HSLS $|E(\mathcal{H})|$ 38 32 32 34
Smoking TV 305.5 431.1 206.1 162.0
Temperature TV 2006.7 922.2 500.3 1184.6